# Entity Framing and Role Portrayal in the News

### v1

@@ Abstract
We introduce a novel multilingual hierarchical corpus annotated for entity framing and role portrayal in news articles. The dataset uses a unique taxonomy inspired by storytelling elements, comprising 22 fine-grained roles, or archetypes, nested within three main categories: protagonist, antagonist, and innocent. Each archetype is carefully defined, capturing nuanced portrayals of entities such as guardian, martyr, and underdog for protagonists; tyrant, deceiver, and bigot for antagonists; and victim, scapegoat, and exploited for innocents. The dataset includes 1,378 recent news articles in five languages (Bulgarian, English, Hindi, European Portuguese, and Russian) focusing on two critical domains of global significance: the Ukraine-Russia War and Climate Change. Over 5,800 entity mentions have been annotated with role labels. This dataset serves as a valuable resource for research into role portrayal and has broader implications for news analysis. We describe the characteristics of the dataset and the annotation process, and we report evaluation results on fine-tuned state-of-the-art multilingual transformers and hierarchical zero-shot learning using LLMs at the level of a document, a paragraph, and a sentence.
@@
Tarek Mahmoud1, Zhuohan Xie1, Dimitar Dimitrov2, Nikolaos Nikolaidis3, Purificação Silvano4, Roman Yangarber5, Shivam Sharma6, Elisa Sartori7, Nicolas Stefanovitch8, Giovanni Da San Martino7 Jakub Piskorski9, Preslav Nakov1, 1MBZUAI, 2Sofia University "St. Kliment Ohridski", 3Athens University of Economics and Business, 4University of Porto, 5University of Helsinki, 6Indian Institute of Technology Delhi, 7University of Padova, 8European Commission Joint Research Centre, 9Institute of Computer Science, Polish Academy of Science {tarek.mahmoud, preslav.nakov}@mbzuai.ac.ae
@@ 1 Introduction
The rapid proliferation of social media has dramatically transformed the information landscape, providing immediate access to news, and allowing anyone to propagate their narratives across the globe. While this connectivity functions as a convenient avenue for information dissemination, it also heightens the risk of exposure to biased reporting, propaganda, and narrative manipulation. These risks are particularly pronounced during periods of conflict and political upheavals, where the framing of entities—individuals, organizations, or groups—can profoundly influence public perception and decision-making. Understanding how entities are portrayed in the news is essential for fostering media literacy, identifying bias, and ensuring transparent news consumption.
Social science highlights the role of emotion in framing—selecting elements that evoke affective responses to shape perceptions Gross and D’Ambrosio (2004). Emotional framing often leverages language that elicits specific feelings, such as fear, anger, or compassion, to influence how entities and events are understood Iyengar (1991); Nabi (2003); Semetko and Valkenburg (2000); Brader (2006). For instance, referring to a group as “freedom fighters” versus “terrorists” not only frames their role but also activates distinct emotional reactions. Research has shown that emotional appeals are powerful tools for reinforcing or challenging public attitudes Westen (2008); Lerner et al. (2015). Such framing can manifest through specific linguistic cues and includes the portrayal of entities also defined by Schneider and Ingram (1993) as the Social Construction of Target Populations.
Natural language processing research has increasingly been applied to analyze the emotional dimensions of framing Troiano et al. (2023), including the identification of sentiment Zhang et al. (2024); Tan et al. (2023) and emotion-laden narratives Mousavi et al. (2022). Understanding these emotional components provides deeper insight into how media narratives construct and perpetuate particular representations of entities, ultimately shaping public perception and societal discourse.
Given the large scale and complexity of the modern news ecosystem, effective analysis of entity framing requires automated tools, which depend on high-quality annotated data. In this context, we introduce a new multilingual dataset designed to develop tools for the study of entity framing and role portrayal in news articles. Our dataset uses a unique, hierarchical taxonomy inspired by elements of storytelling containing a set of 22 carefully defined archetypes nested under three main roles: protagonist, antagonist, and innocent.
@@
  • We provide comprehensive dataset statistics, exploring entity portrayals across languages and topics.
@@ 2 Related Work
Sharma et al. (2023) introduced a dataset for identifying heroes, villains, and victims in memes, focusing on visual content. In contrast, our dataset focuses on textual content. While both share similar coarse-level roles, our work adds an additional layer of granularity with a hierarchical taxonomy of 22 archetypes nested within these roles.
Card et al. (2016) addressed a different aspect of framing. Their contribution is developing a model that makes use of personas to infer the article framing as defined in the Media Frames Corpus (MFC) Card et al. (2015). MFC focuses on identifying how an article is framed along nine dimensions, such as Economic or Political. In their work, topic modeling is used to identify 50 personas. However, the results are noisy, with only a few informative personas, namely refugee and immigrant, while the 48 other personas, such as Job, Worker, and Year, are less informative. In contrast, our work focuses on entity framing rather than article framing. While their work identifies personas in a weakly supervised manner, we develop a hierarchical taxonomy containing a richer set of roles, validated through human annotation of news articles across two diverse domains, ensuring higher quality and broader applicability. There is more research on news framing that focuses on article-level framing Pastorino et al. (2024); Otmakhova et al. (2024); Piskorski et al. (2023); Liu et al. (2019); Card et al. (2015). In contrast, our work centers on entity-level framing.
Aspect-Based Sentiment Analysis Chebolu et al. (2024); Zhang et al. (2022); Orbach et al. (2021); Jiang et al. (2019); Saeidi et al. (2016) is also related. It involves identifying targets of specific opinions and determining the polarity of the sentiment associated with particular aspects of these targets. Typically, the polarity is binary, and multiple aspects of the target entity are examined. Our work on entity framing is different as we do not define aspects nor do we assign polarities. Instead, we introduce a hierarchical taxonomy for news, which contains a rich set of roles inspired by elements of storytelling, and entities can be classified into any subset of roles within that taxonomy.
@@ 3 The Entity Framing Task
Entity framing focuses on analyzing how a text portrays a specific entity through word choice and narrative structure. More concretely, given a news article and a list of entity mentions (i.e., entity mentions, along with their span offsets), we assign to each of them one or more roles based on the taxonomy shown in Figure 2. We developed this taxonomy specifically for this task and the roles were inspired by storytelling elements. The taxonomy includes 22 archetypes, or fine-grained roles nested under three main roles: protagonist, antagonist, and innocent. The role an entity plays in a given article may differ from one context in that article to another depending on the portrayal. See Figure 1 for a complete, annotated example from the corpus.
@@ 4.3 Annotation Process
Given that our corpus contains articles in five languages, we had one annotation team per language, each led by a language coordinator. Each language team included 3 to 5 annotators with prior experience in linguistics, social science, international relations, or prior annotation work. The annotators studied our detailed guidelines, attended live demonstrations, and completed real-time annotation exercises. During weekly meetings, teams clarified any uncertainties, resolved conflicts, improved consistency, and revised the annotation guidelines.
@@
Table 3: Proportion of entities of a given frequency in the corpus.
@@ 4.4 Inter-Annotator Agreement
To assess the inter-annotator agreement (IAA), we used Krippendorff’s alpha. We compared the annotations at the span level: they were approximately matched if they shared at least 50% of their length, to account for minor differences. The IAA values are shown in Table 2. The results indicate a moderate agreement (above 0.45), which we consider acceptable due to the span-based nature of the task. We can see that the IAA is similar across the languages, except for Bulgarian, for which it is notably higher (0.73), which can be explained by the low count of entities in Table 1.
@@ 4.5.1 Statistics
Table 3 presents the number and the proportion of entities within a given frequency range considering exact string matching. We can see that 74% of the entities were annotated only once, while 14 entities have been annotated more than 50 times These numbers consider only the surface string, not accounting for different grammatical and name variants and different languages. In Appendix D, we matched the most frequent entities accounting for these differences and presented detailed statistics at the level of entities.
@@ 4.5.2 Co-occurrence of Roles
In our definition of entity framing, entity mentions can be assigned one or more fine-grained roles. Our corpus contains on average 1.12 annotations per entity mention. Out of the 1,378 articles, 353 articles contain 638 instances where at least one entity mention has multiple annotations. For this set of articles, the average and the maximum are 2.1 and 3 annotations per entity mention, respectively. We further observe that roles such as peacemaker frequently accompany guardian. Similarly, entities portrayed as scapegoats are often framed as being exploited. More details about the co-occurrence matrix for entity mentions with multiple annotations are shown in Figures 5 and 4 in Appendix D.
@@ 5 Experiments
We experimented with classifying entities into main roles and fine-grained roles. As we performed the entity framing annotations at the span level, we framed the problem as a multi-class multi-label classification task. Given an article, an entity mention, and the span offsets, the goal was to predict the framing of that entity mention. We provide two sets of baselines and experiments to benchmark state-of-the-art models, as well as to assess the complexity of the entity framing task. The first set evaluates fine-tuning multilingual transformer models in various settings, while the second set explores hierarchical zero-shot learning using LLMs.
@@ 5.1 Fine-Tuning Pre-trained Multilingual Transformers
For the first set of experiments, we designed our experiments to address the following aspects:
@@ 5.3 Results
Table 5 offers interesting insights into the performance of XLM-R when fine-tuned on monolingual and multilingual data for multi-label fine-grained role classification at the paragraph level. The monolingual setting exhibits varying performance across languages, with the highest scores achieved for Portuguese, while English, Bulgarian, and Russian show notably lower performance; the model’s performance on Hindi is moderate. These differences stem from the quantity and quality of training data, linguistic variations, and the complexity of entity mentions across languages. The consistently low Macro F1 scores across all languages indicate difficulty in predicting rare roles. In contrast, the multilingual setting consistently outperforms the monolingual setting, demonstrating its ability to better capture diverse fine-grained roles through cross-lingual transfer.
@@
Table 6 consolidates the results from fine-tuning XLM-R and hierarchical zero-shot learning, showing that the multi-step approach achieves slightly better performance on main role prediction compared to the single-step approach. This performance improvement in multi-step prompting can be attributed to the structured decomposition of the task. By explicitly isolating the main role prediction into a dedicated step, the model can focus on high-level role identification without the distraction of fine-grained details. While the multi-step approach enhances performance for main role prediction, it underperforms compared to the single-step approach in predicting fine-grained roles. We hypothesize two potential reasons for this discrepancy:
@@
Table 6 additionally compares the performance of zero-shot approaches and XLM-R, both using similar-sized input contexts. We can see that XLM-R outperforms zero-shot methods on all evaluation measures except for Macro F1, where it shows the lowest performance among all approaches. We hypothesize that this discrepancy arises because XLM-R had limited training instances for rare roles, preventing effective learning for these categories. In contrast, zero-shot approaches do not rely on training data and thus are not constrained by this limitation.
@@ 6 Conclusion and Future Work
We presented a novel multilingual and hierarchical dataset for characterizing entity framing and role portrayal in news articles. Our dataset introduces a unique taxonomy inspired by storytelling elements, featuring 22 fine-grained archetypes nested within three main categories: protagonist, antagonist, and innocent. The dataset covers 1,378 recent news articles in five languages (Bulgarian, English, Hindi, European Portuguese, and Russian), spanning two globally significant domains: the Ukraine-Russia War and Climate Change. Over 5,800 entity mentions have been thoroughly annotated with role labels, capturing nuanced portrayals. We evaluated the dataset using fine-tuned state-of-the-art multilingual transformer models and explored hierarchical zero-shot learning with LLMs at document, paragraph, and sentence level. Our experiments highlight the potential of multilingual representations and hierarchical approaches for entity-framing tasks. We intend to release the dataset to the community freely for research purposes. We hope that this dataset will serve as a valuable resource for developing methods and tools to enhance the analysis of entity portrayals in news media.
In future work, we plan to extend the annotations to additional languages and explore other sources of text, such as social media posts. This expansion aims to provide a broader understanding of role portrayal across diverse linguistic and contextual settings. We also intend to perform a more in-depth dataset analysis to examine formulations such as central entity identification. We will also consider evaluating the potential biases in the dataset to ensure robustness in downstream tasks and end-user applications.
7 Limitations
@@
Our dataset focuses on two domains: the Ukraine-Russia War and Climate Change, and covers news articles in five languages (Bulgarian, English, Hindi, Portuguese, and Russian). While these domains and languages provide a diverse foundation for entity framing analysis, the corpus contains 1,378 articles and it should not be considered representative of all news coverage or media landscapes in any specific country. Additionally, the dataset is not perfectly balanced with respect to topics, entities, or languages. Moreover, human annotation of entity framing and role portrayal inevitably is subjective and annotators may subconsciously have biases that could influence the quality. Despite providing detailed annotation guidelines and conducting quality control measures such as double annotation and adjudication through the curation process, some level of subjectivity may remain in the dataset.
@@
8 Ethics and Broader Impact
@@
Our dataset aims to capture a balanced range of perspectives on the Ukraine-Russia War and Climate Change, covering five languages: Bulgarian, English, Hindi, European Portuguese, and Russian. While our goal is to incorporate diverse news sources and viewpoints, achieving perfect balance is not always possible. Consequently, inherent biases in the original media sources may be present in the annotations. To reduce unwanted annotation biases, the corpus is annotated with clear instructions to annotators to focus strictly on the framing of entities, setting aside their personal opinions. All annotations are performed by subject-matter experts, and we did not use crowd-sourcing.

### v2

@@ Abstract
We introduce a novel multilingual hierarchical corpus annotated for entity framing and role portrayal in news articles. [^1] Our dataset is available at
https://mbzuai-nlp.github.io/entity-framing/
The dataset uses a unique taxonomy inspired by storytelling elements, comprising 22 fine-grained roles, or archetypes, nested within three main categories: protagonist, antagonist, and innocent. Each archetype is carefully defined, capturing nuanced portrayals of entities such as guardian, martyr, and underdog for protagonists; tyrant, deceiver, and bigot for antagonists; and victim, scapegoat, and exploited for innocents. The dataset includes 1,378 recent news articles in five languages (Bulgarian, English, Hindi, European Portuguese, and Russian) focusing on two critical domains of global significance: the Ukraine–Russia War and Climate Change. Over 5,800 entity mentions have been annotated with role labels. This dataset serves as a valuable resource for research into role portrayal and has broader implications for news analysis. We describe the characteristics of the dataset and the annotation process, and we report evaluation results on fine-tuned state-of-the-art multilingual transformers and hierarchical zero-shot learning using LLMs at the level of a document, a paragraph, and a sentence.
@@
Tarek Mahmoud1, Zhuohan Xie1, Dimitar Dimitrov2, Nikolaos Nikolaidis3, Purificação Silvano4, Roman Yangarber5, Shivam Sharma6, Elisa Sartori7, Nicolas Stefanovitch8, Giovanni Da San Martino7, Jakub Piskorski9, Preslav Nakov1 1MBZUAI, UAE  2Sofia University "St. Kliment Ohridski", Bulgaria 3Athens University of Economics and Business, Greece  4University of Porto, Portugal 5University of Helsinki, Finland  6Indian Institute of Technology Delhi, India 7University of Padova, Italy  8European Commission Joint Research Centre, Italy 9Institute of Computer Science, Polish Academy of Sciences, Poland {tarek.mahmoud, preslav.nakov}@mbzuai.ac.ae
@@ 1 Introduction
The rapid proliferation of massive amounts of content on social media has dramatically transformed the information landscape, providing immediate access to news, and allowing anyone to propagate their narratives across the globe. While this connectivity functions as a convenient avenue for information dissemination in a very democratic way, it also heightens the risk of online users of being exposed to biased reporting, propaganda, and narrative manipulation.
These risks are particularly pronounced during periods of conflict and political upheavals, where the framing of entities—individuals, organizations, or groups—can profoundly influence public perception and decision-making. Understanding how entities are portrayed in the news is essential for fostering media literacy, identifying bias, and ensuring transparent news consumption.
Social science highlights the role of emotion in framing—selecting elements that evoke affective responses to shape perceptions Gross and D’Ambrosio (2004). Emotional framing often leverages language that elicits specific feelings, such as fear, anger, or compassion, to influence how entities are understood Iyengar (1991); Nabi (2003); Semetko and Valkenburg (2000); Brader (2006). For instance, referring to a group as “freedom fighters” versus “terrorists” not only frames their role but also activates distinct emotional reactions. Research has shown that emotional appeals are powerful tools for reinforcing or challenging public attitudes Westen (2008); Lerner et al. (2015). Such framing can manifest through specific linguistic cues and includes the portrayal of entities also defined by Schneider and Ingram (1993) as the Social Construction of Target Populations.
Natural language processing research has increasingly been applied to analyze the emotional dimensions of framing Troiano et al. (2023), including the identification of sentiment Zhang et al. (2024); Tan et al. (2023) and emotion-laden narratives Mousavi et al. (2022) in text. Understanding these emotional components provides deeper insight into how media narratives construct and perpetuate particular representations of entities, which ultimately shapes public perception and societal discourse.
Given the large scale and complexity of the modern news ecosystem, effective analysis of entity framing requires automated tools, which depend on high-quality annotated data.
In this context, we introduce a new multilingual dataset designed to develop tools for the study of entity framing and role portrayal in news articles. Our dataset uses a unique, hierarchical taxonomy inspired by elements of storytelling containing a set of 22 carefully defined archetypal roles nested under three main frames: protagonist, antagonist, and innocent.
@@
  • We provide comprehensive dataset statistics, and analyze co-occurrence of roles as well as role transitions.
@@ 2 Related Work
2.1 A Hierarchical Taxonomy of Entity Roles
The identification of narrative frames such as heroes, villains, and victims has been studied across various modalities and domains. Sharma et al. (2023) introduced a dataset for identifying heroes, villains, and victims in memes, focusing on visual content. Similarly, Stammbach et al. (2022) demonstrated that LLMs can detect these frames from narrative texts using zero-shot question-answering prompts across diverse domains such as newspaper articles, movie plot summaries, and political speeches. Frermann et al. (2023) integrated elements like conflict and resolution with the framing of key entities as heroes, victims, or villains.
These studies primarily focus on the coarse-grained categorization of entities into the hero-villain-victim frames. Bergstrand and Jasper (2018) analyze this triad and show how these frames are imbued with moral attributes where heroes are shown as benevolent and strong, villains as malevolent and strong, and victims as benevolent and weak. In contrast, our framework redefines these frames based on narrative function, rather than moral judgment. By introducing the frames of protagonist, antagonist, and innocent, we focus on the functional roles entities occupy within a narrative, independent of their moral aspect.
This approach allows for a more nuanced and flexible analysis of entity roles. We further present a hierarchical taxonomy of 22 roles nested within these narrative functional frames. This rich set of roles captures the diverse ways entities are portrayed in textual narratives. Our taxonomy was validated through human annotations across diverse news domains.
2.2 The Entity Framing Task Formulation
Aspect-Based Sentiment Analysis (ABSA)
Prior work on ABSA and entity-level sentiment analysis, which is a special case of ABSA and the most relevant one to our research, explores both the assignment of sentiment to particular attributes of entities and to entities as a whole Saeidi et al. (2016); Jiang et al. (2019); Orbach et al. (2021); Zhang et al. (2023); Chebolu et al. (2024); Kuila and Sarkar (2024); Bastan et al. (2020); Rønningstad et al. (2022). For instance, Tang et al. (2023) presented a dataset for entity-level sentiment classification in financial texts. In contrast, our task formulation differs by not assigning polarities or defining aspects; instead, we classify entities into a rich set of roles inspired by storytelling elements.
Article-level framing
Card et al. (2016) addressed article-level framing by developing a model that utilizes personas to infer framing dimensions defined in the Media Frames Corpus Card et al. (2015). Their topic modeling approach identified 50 personas, but only a few, such as “refugee” and “immigrant,” were found to be informative. Further research on news framing has predominantly focused on article-level analysis. Studies like Liu et al. (2019); Piskorski et al. (2023); Otmakhova et al. (2024); Pastorino et al. (2024) have explored various dimensions of framing in news articles. However, our approach centers on the framing of individual entities within the text, not on article-level framing.
@@ 3 The Entity Framing Task
Entity framing focuses on analyzing how a text portrays a specific entity through word choice and narrative structure. More concretely, given a news article and a list of entity mentions (i.e., entity mentions, along with their span offsets), we assign to each of them one or more roles based on the taxonomy shown in Figure 2. We developed this taxonomy specifically for this task and the roles were inspired by storytelling elements.
The taxonomy includes 22 archetypes, or fine-grained roles nested under three main frames: protagonist, antagonist, and innocent. The role an entity plays in a given article may differ from one context in that article to another depending on the portrayal. See Figure 1 for a complete, annotated example from the corpus.
@@ 4.3 Annotation Process
Given that our corpus contains articles in five languages, we had separate annotation teams, one for each language, each led by a language coordinator. Each language team included 3 to 5 annotators with prior experience in linguistics, social science, international relations, or prior annotation work. The annotators studied our detailed guidelines, attended live demonstrations, and completed real-time annotation exercises. During weekly meetings, teams clarified any uncertainties, resolved conflicts, improved consistency, and revised the annotation guidelines.
@@
Table 3: Proportion of entities of a given frequency.
@@ 4.4 Inter-Annotator Agreement
To assess the inter-annotator agreement (IAA), we used Krippendorff’s alpha. We compared the annotations at the span level: they were approximately matched if they shared at least 50% of their length, to account for minor differences. The IAA values are shown in Table 2. The results indicate a moderate agreement (above 0.45), which we consider acceptable due to the span-based nature of the task.
We can see that the IAA is similar across the languages, except for Bulgarian, for which it is notably higher (0.73), which can be explained by the low count of entities in Table 1.
@@ 4.5.1 Statistics
Figure 3: Top entities’ counts after multilingual linking was manually performed to link surface string to unique identifiers.
@@
Table 3 shows the number and proportion of entities within a given frequency range based on exact string matching. We can see that 74% of the entities were annotated only once, while 14 entities were annotated more than 50 times. These numbers only reflect surface string, not accounting for grammatical and name variants, or languages.
In Figure 3, we match the most frequent entities accounting for these differences and present detailed statistics at the level of entities. The entities selected are all the ones for which at least one surface string has a count of at least 10 in any language.
@@ 4.5.2 Co-occurrence of Roles
In our definition of entity framing, entity mentions can be assigned one or more fine-grained roles. Our corpus contains on average 1.12 annotations per entity mention. Out of the 1,378 articles, 353 articles contain 638 instances where at least one entity mention has multiple annotations. For this set of articles, the average and the maximum are 2.1 and 3 annotations per entity mention, respectively. We further observe that roles such as peacemaker frequently accompany guardian. Similarly, entities portrayed as scapegoats are often framed as being exploited. More details about the co-occurrence matrix for entity mentions with multiple annotations are shown in Figures 5 and 6 in Appendix D.
4.5.3 Analysis of Role Transitions
We analyze role transitions of an entity within the same article, and observed that out of 1,378 articles, only 99 contained occurrences with main role changes for the same entity, and 261 contained fine-grained role changes.
While the most common main role transitions are between protagonist and antagonist roles, fine role transitions are more nuanced. Examples include sequences like victim to underdog to virtuous and transitions from rebel, and virtuous to exploited and victim. We study the transition sequences for both main and fine-grained roles, and present the ten most frequent transitions sequences in Figure 7 in Appendix D.
Our annotation guidelines account for such instances. While such cases are relatively infrequent, they highlight the complexity of entity framing and the importance of capturing role transitions within our taxonomy. This analysis also reinforces the value of our dataset for studying such nuanced challenges.
@@ 5 Experiments
We experimented with classifying entities into main roles and fine-grained roles. As we performed the entity framing annotations at the span level, we framed the problem as a multi-class multi-label classification task. Given an article, an entity mention, and the span offsets, the goal was to predict the framing of that entity mention. We provide two sets of baselines and experiments to benchmark state-of-the-art models, as well as to assess the complexity of the entity framing task. The first set evaluates fine-tuning multilingual transformer models in various settings, while the second set explores hierarchical zero-shot learning using LLMs. To prevent data leakage, we created train/dev/test splits based on entire articles rather than individual entity-mention annotations. The details of these splits are provided in Table 4.
BG EN HI PT RU All
Train 165 (389) 133 (440) 203 (1347) 206 (833) 89 (252) 796 (3261)
Dev 94 (237) 69 (245) 139 (983) 100 (417) 44 (114) 446 (1996)
Test 15 (30) 27 (90) 35 (279) 31 (115) 28 (85) 136 (599)
Total 274 (656) 229 (775) 377 (2609) 337 (1365) 161 (451) 1378 (5856)
@@ 5.1 Fine-Tuning Pre-trained Multilingual Transformers
For the first set of experiments, we designed a series of experiments to address the following important aspects:
@@ 5.3 Results
Table 6 offers interesting insights into the performance of XLM-R when fine-tuned on monolingual and multilingual data for multi-label fine-grained role classification at the paragraph level.
The monolingual setting exhibits varying performance across languages, with the highest scores achieved for Portuguese, while English, Bulgarian, and Russian show notably lower performance; the model’s performance on Hindi is moderate. These differences stem from the quantity and quality of training data, linguistic variations, and the complexity of entity mentions across languages. The consistently low Macro F1 scores across all languages indicate difficulty in predicting rare roles. In contrast, the multilingual setting consistently outperforms the monolingual setting, demonstrating its ability to better capture diverse fine-grained roles through cross-lingual transfer.
@@
Table 7 consolidates the results from fine-tuning XLM-R and hierarchical zero-shot learning, showing that the multi-step approach achieves slightly better performance for main role prediction compared to the single-step approach. This can be attributed to the structured decomposition of the task.
By explicitly isolating the main role prediction into a separate dedicated step, the model can focus on high-level role identification without the distraction of fine-grained details. While this multi-step approach enhances performance for main role prediction, it underperforms compared to the single-step approach for predicting fine-grained roles. We hypothesize two potential reasons for this discrepancy:
@@
Table 7 additionally compares the performance of zero-shot approaches and XLM-R, both using similarly sized input contexts. We can see that XLM-R outperforms zero-shot methods on all evaluation measures except for Macro F1, where it shows the lowest performance among all approaches we experimented with.
We hypothesize that this discrepancy arises because XLM-R had only seen a limited number of training instances for the rare roles, which makes it very hard for the model to learn to detect these categories. In contrast, zero-shot approaches do not rely on training data and thus are not constrained by this limitation.
This hypothesis is based on observable patterns in the dataset and model performance. Figure 4 illustrates the fine-grained role distribution, revealing a significant imbalance, where we can see that certain roles are severely underrepresented. This imbalance intuitively contributes to the observed discrepancy between Micro and Macro F1 scores, as XLM-R’s fine-tuning relies on sufficient training instances for rare roles, which are limited in this dataset. This observation is consistent with previous findings in NLP Henning et al. (2023), where class imbalance is known to affect the performance for low-frequency classes.
In contrast, zero-shot approaches do not depend on labeled training data, which makes them less sensitive to class imbalance, which can explain their comparatively higher Macro F1 scores. While factors such as model architecture or prompt design could also influence performance, dataset imbalance remains a key explanatory factor supported by our analysis. Thus, while XLM-R achieves a higher Micro F1 score, the higher Macro F1 score that we observe for zero-shot approaches offers an intuitive explanation tied to the distribution of roles in the dataset.
@@ 6 Conclusion and Future Work
We presented a novel multilingual and hierarchical dataset for characterizing entity framing and role portrayal in news articles. The dataset introduces a unique taxonomy inspired by storytelling elements, featuring 22 fine-grained archetypes nested within three main categories: protagonist, antagonist, and innocent. The dataset covers 1,378 recent news articles in five languages (Bulgarian, English, Hindi, European Portuguese, and Russian), spanning two domains: the Ukraine–Russia War and Climate Change. Over 5,800 entity mentions have been thoroughly annotated with role labels, capturing nuanced portrayals. We evaluated the dataset using fine-tuned state-of-the-art multilingual transformers and explored hierarchical zero-shot learning at the document, paragraph, and sentence level. Our experiments highlighted the potential of multilingual representations and hierarchical approaches for entity-framing tasks. We release the dataset to the community freely for research purposes, and we hope it will enable future research.
In future work, we plan to extend the annotations to additional languages and explore other sources, such as social media posts. This expansion aims to provide a broader understanding of role portrayal across diverse linguistic and contextual settings. We also intend to perform a more in-depth dataset analysis to examine formulations such as central entity identification. We further plan to evaluate the potential biases in the dataset aiming to ensure robustness for downstream tasks and end-user applications.
Limitations
@@
Our dataset focuses on two domains: the Ukraine–Russia War and Climate Change, and covers news articles in five languages (Bulgarian, English, Hindi, Portuguese, and Russian). While these domains and languages provide a diverse foundation for entity framing analysis, the corpus contains 1,378 articles and it should not be considered representative of all news coverage or media landscapes in any specific country. Moreover, the dataset is not balanced with respect to topics, entities, or languages. Furthermore, human annotation of entity framing and role portrayal inevitably is subjective and annotators may subconsciously have biases that could influence quality. Despite providing detailed annotation guidelines and conducting quality control measures such as double annotation and adjudication through the curation process, some level of subjectivity may remain in the dataset.
@@
Ethics and Broader Impact
@@
Our dataset aims to capture a balanced range of perspectives. While our goal is to incorporate diverse news sources and viewpoints, achieving perfect balance is not always possible. Consequently, inherent biases in the original media sources may be present in the annotations. To reduce unwanted annotation biases, the corpus is annotated with clear instructions to annotators to focus strictly on the framing of entities, setting aside their personal opinions. All annotations are performed by subject-matter experts, and we did not use crowd-sourcing.
@@
Acknowledgments
This research is partially funded by the EU NextGenerationEU, through the National Recovery and Resilience Plan of the Republic of Bulgaria, project SUMMIT, No BG-RRP-2.004-0008.
Giovanni Da San Martino is funded by the European Union under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 - Call for tender No. 341 of March 15, 2022 of Italian Ministry of University and Research – NextGenerationEU; Code PE00000014, Concession Decree No. 1556 of October 11, 2022 CUP D43C22003050001, Progetto SEcurity and RIghts in the CyberSpace (SERICS) - Spoke 2 Misinformation and Fakes - DEcision supporT systEm foR cybeR intelligENCE (Deterrence).